Hello, I have been working on this problem all day and it seems the more I work on it the more confusing I seem to make it for myself. I over think too much if that is even possible Here is the problem: There are 850 Douglas fir and ponderosa pine trees in a section of forest bought by Sawz Logging Co. The company paid an average of $300 for each Douglas fir and $225 for each ponderosa pine. If the company paid $217,500 for the trees, how many of each kind did the company buy?
I would really appreciate your help here I'm sure its so simple but I'm just not getting it.
Thank You, Barbara Neely

Answer by checkley71(8403)

D+P=850 or D=850-P NOW SUBSTITUTE (850-P) FOR D IN THE FOLLOWING EQUATION & SOLVE FOR P
300D+225P=217,500
300(850-P)+225P=217,500
255,000-300P+225P=217,500 [THE 255,000 IS THE PRODUCT OF (300*850)]
-75P=217,500-255,000
-75P=-37,500
P=-37,500/-75
P=500 IS THE NUMBER OF PONDEROSA PINES BOUGHT.
850-500=350 IS THE NUMBER OF DOUGLAS FIRS BOUGHT.
PROOF
300*350+225*500=217,500
105,000+112,500=217,500
217,500=217,500